# Optimal Quoting under Adverse Selection and Price Reading

### v1

@@ 5.3 Adverse Selection
Let us now consider the case of adverse selection only. We have
@@
As discussed above, the first term represents the global response to adverse selection risk, for trades of size $\Delta^{k}$, uniformly across all tiers, whereas the second term is specific to each tier. The magnitude and sign of these effects depend critically on the relative slopes of the exponential functions defining the intensity and price impact functions – that is, on the differences $\beta^{m,j}-\kappa^{m,j}$ appearing in the first sum, and on the tier-specific difference $\beta^{n,k}-\kappa^{n,k}$ in the second term.
As an illustration, we consider two tiers with equal intensities, as in the previous example. The first tier is assumed to be standard, with $\alpha^{1,k}=0$, while the second tier is endowed with a size-uniform informational advantage, characterized by $\alpha^{2,k}=0.1$, and slow signals, i.e., $\beta^{2,k}<\kappa^{2,k}$, here with $\frac{\beta^{2,k}}{\kappa^{2,k}}=0.9$.
Figure 3 shows that, in the presence of such adverse selection, the market maker behaves as if they were less risk averse when it comes to the quotes shown to the first tier. In particular, they tighten the bid-ask spread for the first tier when inventory is zero. By contrast, the spread quoted to the second (informed) tier is widened. Nevertheless, the market maker can still exploit the top-of-book prices shown to the informed tier to probe price direction. The limited effect on larger-size quotes is partly due to our modelling assumptions, but the underlying mechanism is general: small risk-reducing trades with informed clients cause limited losses while increasing the value of the remaining inventory.
A key takeaway from this analysis is that an informed tier subject to adverse selection should not necessarily be avoided unlike what initial intuition could suggest. Rather, it can be leveraged as a form of “signal subscription,” beneficial for the risk management of the entire franchise.
@@
Figure 3: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$, for regular tier and tier with adverse selection. Risk aversion $\gamma=10^{-5}$ bp-1 M-1, other parameters are defined in the text.
@@ Conclusion
In this article, we have proposed a tractable and implementable stochastic optimal control framework that enables market makers to adjust their quotes with greater awareness of informational risk arising from both price reading and adverse selection.
Using perturbation analysis, we have found that the impact of informational risk can be split into two components: (i) a global component related to induced changes of the value function that applies to pricing for all clients, and (ii) an individual component that applies to specific tiers to mitigate their perceived informational risk.
In the presence of price reading, the market maker acts as if they were more risk averse when showing prices to their global franchise. At the same time, they reduce the informational content of their quotes shown to skew sniffers. The extent of skew reduction strongly depends on liquidity, specifically on the contribution of price readers to market maker’s risk reduction. In particular, it is not beneficial to show skew to price readers with little liquidity contribution.
Optimal response to adverse selection depends on the relative shape of informed flow intensity and price impact functions. The market maker is able to become less risk averse for their global franchise when informed clients trade on long-term signals. Top-of-book prices for risk-reducing trades will be very attractive for informed clients. This is the cost the market maker is willing to pay to probe price direction for the benefit of a better risk management of their overall portfolio. On the contrary, adverse selection with sharp signals forces the market marker to be protective by showing defensive quotes even for risk-reducing trades with informed clients and behave as if they were more risk averse when it comes to their global franchise.
@@ Acknowledgment
We would like to express our sincere gratitude to all those who made possible the research that led to this paper. The authors are especially grateful to Richard Anthony (HSBC).
@@
The data that support the findings of this article are available from the authors upon reasonable request.

### v2

@@ 5.3 Adverse Selection
Let us now consider the case of adverse selection alone. We have
@@
As discussed above, the first term represents the global response to adverse selection risk, for trades of size $\Delta^{k}$, uniformly across all tiers, whereas the second term is tier-specific. The magnitude and sign of the different effects depend critically on the relative slopes of the exponential functions defining the intensity and price impact functions – that is, on the differences $\beta^{m,j}-\kappa^{m,j}$ appearing in the first sum, and on the tier-specific difference $\beta^{n,k}-\kappa^{n,k}$ in the second term.
As an illustration, we consider the same market as in Section 5.2 with a similar two-tier set up. The first tier is assumed to be standard, with $\alpha^{1,k}=0$, while the second tier is endowed with a size-uniform informational advantage, characterized by $\alpha^{2,k}=0.05$, and slow signals, i.e., $\beta^{2,k}<\kappa^{2,k}$, here with $\beta^{2,k}=2$.
Figure 6: Optimal bid quote adjustment, $\varepsilon^{-1}\left(\widehat{\mathfrak{d}}^{n,k,b*}(q)-\widehat{\delta}^{n,k,b*}(q)\right)$, with the second tier subject to adverse selection (we illustrate the first and the fourth sizes) – $\text{SVS}=50\%$.
Figure 6 illustrates the impact of adverse selection on quotes when $\text{SVS}=50\%$. We see that, in the presence of adverse selection, the market maker behaves as if they were less risk averse when it comes to the quotes shown to the first tier. In particular, they tighten the bid-ask spread for the first tier when inventory is zero. By contrast, the spread quoted to the second (informed) tier is widened. Nevertheless, the market maker can still exploit the top-of-book prices shown to the informed tier to probe price direction. The limited effect on larger-size quotes is partly due to our modelling assumptions, but the underlying mechanism is general: small risk-reducing trades with informed clients cause limited losses while providing meaningful information. A key takeaway from this analysis is that a tier subject to adverse selection should not necessarily be avoided by the market maker unlike what initial intuition could suggest. Rather, it can be leveraged as a form of “signal subscription,” beneficial for the risk management of the entire franchise.
@@
Let us now simulate the PnL of a market maker facing adverse selection as described above, over a time window $[0,T]$ where $T=10^{4}$ s, for different values of the safe volume share SVS and two quoting strategies: (i) the “No Action” strategy, which does not modify quotes to take adverse selection into account and therefore uses the quotes from Eq. (11), and (ii) the “Optimal” strategy, which uses the first-order approximations of Eqs. (12) and (13) with $\varepsilon=1$.
Interestingly, unlike in the case of price reading, there is no obvious naïve strategy to use for comparison in the case of adverse selection. The only alternative would be to stop pricing to the suspected tier, which would drastically damage the PnL and make any comparison irrelevant.
Figures 7, 8, and 9 illustrate the average PnL, its standard deviation, and the ratio between the two (as a risk-adjusted performance measure), respectively, over $10^{5}$ independent simulations for several values of SVS when using the above two strategies. The “No Adverse Selection” baseline corresponds to the absence of adverse selection, i.e., $\alpha^{2,k}=0$, and uses the quoting strategy of Eq. (11). We clearly see that adverse selection negatively impacts the expected PnL while the risk remains almost identical (up to a slight decrease) when no action is taken. The use of our first-order approximations of the optimal pricing strategy helps to protect the expected PnL and significantly reduces risk, particularly when the share of informed clients is moderate. The risk-adjusted performance measures in Figure 9 clearly emphasize the benefits of our first-order approximations, with no clear dependence on SVS.
Figure 7: Average values of the simulated market maker’s PnL as a function of safe volume share, comparing the “Optimal” pricing strategy with the baseline case of no adverse selection and adverse selection with no strategic action by the market maker.
Figure 8: Values of the standard deviation for the simulated market maker’s PnL as a function of safe volume share, comparing the “Optimal” pricing strategy with the baseline case of no adverse selection and adverse selection with no strategic action by the market maker.
Figure 9: Risk-adjusted performance measure associated with for the simulated market maker’s PnL as a function of safe volume share, comparing the “Optimal” pricing strategy with the baseline case of no adverse selection and adverse selection with no strategic action by the market maker.
@@ Conclusion
In this article, we have proposed a tractable stochastic optimal control framework that enables market makers to adjust their quotes to better account for informational risk arising from both price reading and adverse selection.
Using perturbation analysis, we have found that the impact of informational risk can be split into two components: (i) a global component related to induced changes in the value function that applies to pricing for all clients, and (ii) a tier-specific component to mitigate the perceived informational risk associated with each tier.
In the presence of price reading, the market maker acts as if they were more risk averse when showing prices to their global franchise. At the same time, they reduce the informational content of their quotes shown to skew sniffers. The extent of skew reduction strongly depends on liquidity, specifically on the contribution of price readers to the market maker’s risk reduction.
Optimal response to adverse selection depends on the relative shape of informed flow intensity and price impact functions. The market maker can become less risk averse for their global franchise when informed clients trade on long-term signals. Top-of-book prices for risk-reducing trades will then become very attractive for informed clients. This is the cost the market maker is willing to pay to gain information on price direction, thereby improving the risk management of their overall portfolio. In contrast, adverse selection with sharp signals forces the market maker to act protectively by showing defensive quotes even for risk-reducing trades with informed clients and to behave as if they were more risk averse with their global franchise.
@@ Acknowledgment
The authors would like to express their sincere gratitude to all those who made this research possible. They are especially grateful to Richard Anthony (HSBC) and the late Jean-Michel Beacco (Institut Louis Bachelier), who tragically died while we were writing this paper.
@@
Statement on AI Assistance
The authors acknowledge the use of generative AI language models (ChatGPT and Gemini) to assist with language editing, proofreading, and improving the clarity of this article. These tools did not contribute to the original ideas, analysis, or conclusions. All intellectual content is the authors’ own, and they bear full responsibility for the final text, including its accuracy.
@@
The data supporting the findings of this study were generated via numerical computations and simulations. The models, methods, and parameters used to produce these data are fully described within the manuscript, enabling the reproduction of the results. The code used for this study is not publicly deposited. Reasonable inquiries regarding the replication of the findings may be directed to the authors via email.